OpenStudy (anonymous):

find dy/dx y=a^3+(cos^3)(x)

OpenStudy (lalaly):

did u try to do it?

OpenStudy (anonymous):

Hint let u = cos(x) and do the chain rule. That's assuming a is a constant :-)

OpenStudy (lalaly):

\[\large{\frac{d}{dx}\cos^nx=-nsin(x)\cos^{n-1}(x)}\]

OpenStudy (anonymous):

thanks for the rule , and the hint. that was helpful

OpenStudy (lalaly):

:)